OpenStudy (anonymous):

Given the equation 4x + 2y = 9, which equation below would cause an inconsistent-independent system? 6x + 12y = 13 2x + y = -6 x - 2y = 7 8x - 4y = -5

OpenStudy (anonymous):

MEDAL AND FAN

OpenStudy (anonymous):

@demonchild99

OpenStudy (anonymous):

@Jackeline1928

OpenStudy (anonymous):

@TheSmartOne @texaschic101 @tanya123 @Teddyiswatshecallsme

OpenStudy (anonymous):

Anyone please

OpenStudy (anonymous):

Texaschic

OpenStudy (texaschic101):

Inconsistent means there is no solution...the lines will be parallel. Parallel lines will have the same slope. 4x + 2y = 9 2y = -4x + 9 y = -2x + 9/2 slope here is -2, so we need the other equation to have a slope of -2. 2x + y = -6 y = -2x + 6 slope here is -2 as well. And since they have the same slopes and different y intercepts, this makes them parallel, making them inconsistent. I do not see how a set of equations can be inconsistent and independent because inconsistent means there is no solutions and independent means there is 1 solution.